OpenStudy (anonymous):

4x + 20 = 1/2 * (3x + 5) How do i solve this?

OpenStudy (anonymous):

\[4x+20=\frac{ 1 }{ 2 } \times (3x+5)\]\[4x+20=\frac{ 3x }{ 2} + \frac{ 5 }{ 2 }\]\[8x+40=3x+5\]\[5x=-35\]\[x=-7\]

OpenStudy (anonymous):

Thank you SO much!!! @twvogels !!!

OpenStudy (anonymous):

Also: You don't really need step 2, it just makes it easier to read. You can just multiply everything by 2 and go from Step 1 to Step 3. I just find it easier to read if I show every step.

OpenStudy (anonymous):

okay thx!